# Empowering Visual Artists with Tokenized Digital Assets with NFTs

### v1

@@ 4.2 NFT Overview
4.2.1 Deployment Deployers and Users
In the overall landscape of NFTs, two key roles emerge: the deployer and the user. The deployer, often an artist, creator, or technologist, wields the ability to manifest digital assets into the blockchain realm. As the central deployer in this process, they orchestrate the creation and deployment of NFT smart contracts, sculpting the essence of token standards, metadata intricacies, and additional functionalities. They usually have the admin permission for the NFT Smart Contracts[206], which means they control the particularities of NFT Smart Contracts. With each deployment, the artwork is endowed a new digital ownership. The other key role is maintained by users, including collectors, investors, or enthusiasts. As expected, they browse NFT marketplaces and platforms to discover digital art, collectibles, and virtual assets, interacting with NFT smart contracts with normal permissions. Through their activities, they give these NFT value and create new stories about ownership and cultural importance, which are often shared or even celebrated on other social media platforms such as LinkedIn, TikTok, Facebook, and X (formerly Twitter).
4.2.2 NFT Structure
NFTs have a comprehensive structure that defines how they work (cf. Figure 4). Each NFT belongs to a collection and has a unique identifier in the collection called a token ID, which differentiates it from others. Every token ID has its own metadata, which includes important properties like the name, description, creator, image, and other relevant information about the "minted" digital asset This metadata helps explain the NFT’s background detailed.
@@ 4.2.2 NFT Structure
Furthermore, NFTs often use off-chain storage to store metadata and digital asset. There is a link between NFT and metadata called tokenURI, which is a standard way to link to the metadata and digital asset information. The tokenURI usually includes a URL that points to where the metadata are stored, which makes it easy for users and apps to access and show NFT’s information. Some developers save the encoded metadata with Base64 into tokenURI. Users need to view information after decoding with Base64.
4.2.3 Metadata
NFT metadata is at the core of defining the characteristics of NFTs[18]. NFTs represent unique digital assets on the blockchain network that are verified and tradable, with their token ID and information primarily determined by metadata. The token ID ensures that each token is unique and identifiable, while the metadata describes the attributes and characteristics of the NFT. Although metadata is usually stored off-chain, it is linked to the NFT via the tokenURI on the blockchain. Metadata includes various attributes such as the title, description, image URI, creator information, and other attributes. The title provides the name of the NFT and artwork, while the description explains what the NFT represents in terms of the artwork. The image URI links to the digital art or digital assets stored in other storage. Metadata is typically formatted in JSON[154], which is easy to read and parse. This structured format is crucial not only for human readability but also for machine compatibility. For storage services, centralized storage servers can be used, but decentralized storage solutions like InterPlanetary File System (IPFS)[20] are preferred due to their resilience and resistance to censorship.
@@ 4.3 NFT Lifecycle
The lifecycle of an NFT involves several stages[225]. Figure 5 illustrates their processes.
@@
Creation. This is where the NFT is created or minted by an artist or other creator. During this phase, which has five steps, the digital asset is tokenized on a blockchain network, using smart contracts to ensure its uniqueness and ownership status.
@@
Ownership. The buyer now owns the NFT and can display it in their digital wallet or showcase it in virtual environments, depending on the type of asset. Ownership is verified through cryptographic keys associated with the buyer’s wallet address.
@@
Destruction. In some cases, creators may choose to retire or decommission their NFTs, removing them from circulation and potentially increasing the rarity of existing tokens. This can happen for various reasons, such as the end of a limited edition series or changes in the creator’s preferences. Deployer or users can destroy their NFT by the burn method with their NFT tokenId. However, some NFT is build with permission control, the burn method might be limited to use for users.
@@ 4.4 NFT Standards
Batch Transfers
Easily Query Owner
Support semi-fungible tokens
Interoperability with ERC-20
@@
Use cases Artwork, CryptoKitties [40], BAYC [230] Collectibles, Azuki [16] Enjin [57], Gaming, Artwork Solv Protocol [163] Pandora [147]
The Table 5 summarizes the properties of the NFT standards.
@@
ERC-721A. ERC-721A[90] is a proposed extension of the ERC-721 standard. The purpose of ERC-721A is to enhance the functionality and usability of NFTs on the Ethereum blockchain. ERC-721A is proposed by James Sangalli and the Ethereum community. Moreover, ERC-721A has additional functions to ERC-721, such as batch minting, which executes the creation of multiple tokens in a single transaction. The purpose of ERC-721 is to improve efficiency and reducing gas costs.
ERC-2981: NFT Royalty Standard. ERC-2981[231] is a proposed Ethereum Improvement Proposal (EIP), also known as "EIP-2981", which introduces a standardized interface for royalties on NFTs. ERC-2981 is proposed by William Entriken and David Moore in 2021. ERC-2981 helps to compensates creators and original rights holders for their artwork in the NFT ecosystem. This standard imports a new feature that allows for the automatic and transparent distribution of royalties to the original creators whenever NFTs are resold on the secondary markets.
@@
Now we advance the discussion by focusing on the actual process involved in making an NFT.
@@ 5.1 Procedures for NFT Minting and Trading in the Marketplace
  • Mint. User can mint their NFT from the project website or NFT marketplace. Some NFT marketplace have friendly tools to help creators deploy a new NFT.
@@
  • Ownership Transfer. When someone purchases an NFT, the ownership of that digital asset is transferred to their digital wallet. The transaction is recorded on the blockchain, providing a transparent and immutable record of ownership.
@@ 5.2.1 Marketplaces
Blur [24] is an innovative NFT marketplace that caters to professional traders. Launched in 2022, Blur quickly surpassed OpenSea in terms of volume due to its low fees, high speed, and unique features such as NFT floor depth charts. Blur is more than just a marketplace; it’s an ecosystem consisting of an NFT marketplace, aggregator[39], and lending platform[47]. The Blur marketplace provides advanced analytics, portfolio management features, and the ability to compare NFTs on different marketplaces. The marketplace’s main feature is that it allows for gasless transactions, which significantly reduces transaction costs for users[161]. To get started on Blur, users first connect their wallet and sign a gasless transaction. Then, they can start listing and bidding on NFTs. The platform also encourages users to upload or list NFTs to qualify for airdrops. Blur has its native token, BLUR, which has been distributed to the community over several quarters.
@@
SuperRare [179] is a leading platform for buying, selling, and collecting digital art as NFTs. Launched in 2018, it focuses on high-quality, limited-edition digital artwork created by both emerging and established artists. One of SuperRare’s key features is its curation process, which ensures that top-notch artwork is featured on the platform. Artists must apply or be invited to join SuperRare, and each piece of art goes through a thorough review by the platform’s curation team. This strict curation helps maintain the quality and integrity of the art on SuperRare, making it a trusted marketplace for collectors. The platform offers a vibrant marketplace where users can discover and acquire digital art in various styles, media, and themes. SuperRare supports different formats, including static images, animated GIFs, and even interactive 3D experiences, allowing artists to explore new ways of expressing themselves digitally. A unique aspect of SuperRare is its focus on provenance and royalties. Every time an artwork is resold on the platform, a portion of the sale proceeds automatically goes to the original creator, ensuring that artists continue to benefit from the increasing value of their work over time.
@@
Magic Eden [124] is a decentralized marketplace and ecosystem based on the Solana blockchain, focusing on buying, selling, and trading NFTs. Launched to meet the growing demand for NFTs and digital collectibles, Magic Eden offers a range of features designed to provide a smooth experience for users within the Solana ecosystem. At its core, Magic Eden is a marketplace where users can discover, buy, and sell various NFTs, including digital art, collectible cards, virtual items, and more. The platform has an intuitive interface that makes it easy to browse and explore different collections, helping users find unique and valuable NFTs that match their interests. One of Magic Eden’s main advantages is its integration with the Solana blockchain, known for its high performance, scalability, and low transaction fees. By leveraging Solana’s capabilities, Magic Eden provides fast and cost-effective transactions, allowing users to buy, sell, and trade NFTs with minimal friction. Additionally, Magic Eden offers features like staking, governance, launchpads, and community engagement. Users can stake SOL tokens to earn rewards, participate in governance decisions, and contribute to the platform’s growth.
The Table 6 shows the main features for the NFT marketplace. The Figure 7 provides the NFT marketplace share from Dec 2021 to Sep 2024.
5.2.2 NFT Pricing
@@
Title Data Source Time Period Topic Category AI used AI Model Analysis Data Analysis
@@
NFTs represent a unique class of digital assets, often associated with digital art, collectibles, and virtual real estate. Unlike fungible cryptocurrencies, each NFT is distinct, often carrying unique metadata that determines its individual value. Moreover, there is much research on NFT prices, as shown in Table 7. Understanding the price determinants of NFTs involves different factors that includes textual and visual data, market dynamics, social media influence, and broader financial trends. Textual descriptions and image data within NFT collections can help explain price variations among individual NFTs; however, features extracted from these data do not generalize well to new, unseen collections[237]. This indicates that, while certain stylistic or descriptive elements might boost an NFT’s value within a specific collection, these factors do not necessarily apply across different collections. Interestingly, NFTs described with negative sentiment tend to fetch higher prices than those with positive sentiment[146]. This could be attributed to the perceived uniqueness or controversial nature of such descriptions, which might attract more attention and demand.
@@
NFT markets are relatively immature and often inefficient[52]. For instance, the pricing of virtual real estate in Decentraland does not always reflect all available information, though a general rise in value has been observed over time. This inefficiency suggests opportunities for speculative gains but also underscores the risks involved. NFTs have been found to offer significant diversification benefits within traditional asset-based portfolios[107]. During stable market times, NFTs can act as transmitters of systemic risk, but in the stressful periods, they tend to absorb risk spillovers[9]. This dual role highlights their potential as both a hedge and a diversification tool in investment portfolios. The pricing and popularity of NFTs are heavily influenced by the performance of major cryptocurrencies like Bitcoin and Ether[158]. Increases in cryptocurrency prices could drive hype and attention towards NFTs, as evidenced by co-movement patterns observed through wavelet coherence analysis[53]. Furthermore, Bitcoin returns have been shown to predict NFT growth in popularity, as measured by Google search queries.
The key factors contributing to NFT value have been identified: rarity, reputation, authenticity, significance, utility, interoperability, demand, provenance, royalties, and hype[115]. These aspects encapsulate both intrinsic qualities of NFTs and external market perceptions that drive their valuation. During extreme market conditions, various economic indicators such as equity and gold market uncertainties, business conditions, and geopolitical risks significantly predict returns in specific NFT markets[193]. This indicates that broader economic factors can play a crucial role in determining NFT prices. The valuation of NFTs is influenced by a complex interplay of textual and visual data, social media presence, market dynamics, financial contagion, and broader economic conditions. These diverse factors collectively shape the pricing patterns and perceived value of NFTs, making them a unique and evolving asset class within the digital economy.
@@ 5.3 NFT Security
The rapid proliferation of NFTs market is facing with challenges in cybersecurity that pose significant risks to artists and collectors alike, encompassing both human and technological issues. Figure 8 shows most of the security issues for NFT trading.
@@
Technological vulnerabilities are common because of the complex nature of smart contracts and the intricate design of NFTs. Smart contracts, which are self-executing contracts with the terms of the agreement directly written into code[101], are foundation to the operation of NFTs. However, they are not immune to security flaws. Identity verification issues might allow malicious attackers to impersonate legitimate users and lead to unauthorized transactions. The verifiability of NFT contracts is another critical issue. NFT contract is authentic and has not been tampered with is essential for maintaining trust in the system. Smart contract vulnerabilities, such as reentrancy, transaction order dependence, integer overflows, and unhandled exceptions, all highlight the shared security challenges between NFTs and broader DeFi ecosystems[26, 68, 76, 92, 97, 122].
@@
In addition to these technological issues, the NFT market suffers from a lack of transparency[38], particularly in off-chain trades where transactions occur outside the blockchain. This opacity can facilitate fraudulent activities and make it difficult to verify the authenticity and ownership history of an NFT. Fairness in bidding processes is another concern, with practices such as shill bidding and bid shielding undermining the integrity of auctions. Royalty and marketplace fee evasion pose financial risks to artists and platforms, while cross-platform issues and non-enforcement of contractual terms can further exacerbate financial risks.
Traditional Web2 security issues, such as Denial of Service (DoS) attacks[139], also persist in the NFT ecosystem, threatening the availability and reliability of NFT marketplace platforms. With substantial funds flowing into DeFi applications, the allure of scams and cyber attacks has grown. Cryptoeconomic strategies, such as transaction reordering[100, 118], flash loan abuse[164], arbitrage opportunities[44], and pump-and-dump schemes[221, 98], have been adapted to exploit vulnerabilities in the NFT market.
@@ 5.3.1 Wash Trading
The emergence of NFTs has ushered in a new era of digital asset markets, characterized by unprecedented innovation and rapid growth. However, amidst the excitement surrounding NFTs lies a darker underbelly: the prevalence of wash trading (cf. Figure 9). Wash trading, a deceptive practice involving the simultaneous buying and selling of assets to create artificial trading volume, has raised significant concerns within the NFT ecosystem. Recent studies challenged previous estimates, suggesting that wash trading may be less common than initially thought[194]. Nonetheless, findings still indicate alarming levels of wash trading activity, with some estimates it attributing up to 25% of total trading volume to this illicit practice[189].
@@
Identifying and combating wash trading in NFT markets presents unique challenges. Innovative solutions are being proposed, such as algorithms to detect wash trading at different levels, from transaction to collector[182]. These efforts aim to provide regulators and investors with tools to navigate the complexities of the NFT landscape and mitigate the risks associated with market manipulation. Moreover, the motivations behind wash trading are multifaceted. While some engage in wash trading to artificially inflate NFT prices, others exploit token reward systems offered by marketplaces [113]. The profitability of wash trading is evident, with some perpetrators avoiding significant losses and even profiting from reselling NFTs to unsuspecting buyers.As the NFT market continues to evolve, the need for robust regulation becomes increasingly apparent. Measures must be implemented to prevent market abuse, instill investor confidence, and safeguard the integrity of digital asset markets. Novel visualization tools, like NFTDisk[210], offer investors a means to identify wash trading activities and make informed decisions in an increasingly complex market landscape. The Table 8 provides the summary of the identified literature for NFT Wash Trading.
@@
Unveiling the Paradox of NFT Prosperity[83] OpenSea, X2Y2, CryptoPunks, LooksRare, Blur ERC-721 ERC-1155 Wash trading analysis ACM
@@
Identifying Washtrading Cases in NFT Sales Networks[182] Nifty Gateway Wash trading detection method IEEE Transactions
@@
Beyond the Surface: Advanced Wash Trading Detection in Decentralized NFT Markets[189] Blur, OpenSea, X2Y2, LooksRare ERC-721 Wash trading analysis arVix
@@
NFTDisk: Visual Detection of Wash Trading in NFT Markets[210] OpenSea, LooksRare ERC-721 Wash trading detection tool ACM

### v2

@@ 4.2 NFT Overview
Deployers and Users. In the overall landscape of NFTs, two key roles emerge: the deployer and the user. The deployer, often an artist, creator, or technologist, wields the ability to manifest digital assets into the blockchain realm. As the central deployer in this process, they orchestrate the creation and deployment of NFT smart contracts, sculpting the essence of token standards, metadata intricacies, and additional functionalities. They usually have the admin permission for the NFT Smart Contracts [206], which means they control the particularities of NFT Smart Contracts. With each deployment, the artwork is endowed a new digital ownership. The other key role is maintained by users, including collectors, investors, or enthusiasts. As expected, they browse NFT marketplaces and platforms to discover digital art, collectibles, and virtual assets, interacting with NFT smart contracts with normal permissions. Through their activities, they give these NFT value and create new stories about ownership and cultural importance, which are often shared or even celebrated on other social media platforms such as LinkedIn, TikTok, Facebook, and X (formerly Twitter).
NFT Structure. NFTs have a comprehensive structure that defines how they work (cf. Figure 4). Each NFT belongs to a collection and has a unique identifier in the collection called a token ID, which differentiates it from others. Every token ID has its own metadata, which includes important properties like the name, description, creator, image, and other relevant information about the "minted" digital asset This metadata helps explain the NFT’s background in detail.
Furthermore, NFTs often use off-chain storage to store metadata and digital assets. There is a link between NFT and metadata called tokenURI, which is a standard way to link to the metadata and digital asset information. The tokenURI usually includes a URL that points to where the metadata are stored, which makes it easy for users and apps to access and show NFT’s information. Some developers save the encoded metadata with Base64 into tokenURI. Users need to view information after decoding with Base64.
@@
Metadata. NFT metadata is at the core of defining the characteristics of NFTs [18]. NFTs represent unique digital assets on the blockchain network that are verified and tradable, with their token ID and information primarily determined by metadata. The token ID ensures that each token is unique and identifiable, while the metadata describes the attributes and characteristics of the NFT. Although metadata is usually stored off-chain, it is linked to the NFT via the tokenURI on the blockchain. Metadata includes various attributes such as the title, description, image URI, creator information, and other attributes. The title provides the name of the NFT and artwork, while the description explains what the NFT represents in terms of the artwork. The image URI links to the digital art or digital assets stored in other storage. Metadata is typically formatted in JSON [154], which is easy to read and parse. This structured format is crucial not only for human readability but also for machine compatibility. For storage services, centralized storage servers can be used, but decentralized storage solutions like InterPlanetary File System (IPFS) [20] are preferred due to their resilience and resistance to censorship.
@@ 4.3 NFT Lifecycle
The lifecycle of an NFT involves several stages [225] (cf. Figure 5).
@@
Creation. During this phase, the digital asset is tokenized on a blockchain network, using smart contracts to ensure its uniqueness and ownership status.
@@
Ownership. The buyer owns NFT and can display it in their digital wallets or showcase it in virtual environments, depending on the type of asset. Ownership is verified through cryptographic keys associated with the buyer’s wallet address.
@@
Destruction. In some cases, creators may choose to retire their NFTs, removing them from circulation to increase the rarity of existing tokens. This can happen for various reasons, such as the end of a limited edition series or changes in the creator’s preferences. Deployer or users can destroy their NFT by the burn method with their NFT tokenId. However, some NFT is build with permission control, the burn method might be limited to use for users.
@@ 4.4 NFT Standards
Batch Transfer?
Easily Query Owner?
Support semi-fungible?
Compatible (ERC20)?
@@
Use cases Artwork, CryptoKitties [40], BAYC [230] Collectibles, Azuki [16] Enjin [57], Gaming, Artwork Solv [163] Pandora [147]
@@
ERC-721A. ERC-721A [90] is a proposed extension of the ERC-721 standard. The purpose of ERC-721A is to enhance the functionality and usability of NFTs on the Ethereum blockchain. ERC-721A is proposed by James Sangalli and the Ethereum community. Moreover, ERC-721A has additional functions to ERC-721, such as batch minting, which executes the creation of multiple tokens in a single transaction. The purpose is to improve efficiency and reducing gas costs.
ERC-2981: NFT Royalty Standard. ERC-2981 [231] is a proposed Ethereum Improvement Proposal (EIP), also known as "EIP-2981", which introduces a standardized interface for royalties on NFTs. ERC-2981 is proposed by William Entriken and David Moore in 2021. ERC-2981 helps to compensates creators and original rights holders for their artwork in the NFT ecosystem. This standard imports a new feature that allows for the transparent distribution of royalties to the original creators whenever NFTs are resold on the secondary markets.
@@
Then, we focuse on the actual process in making an NFT.
@@ 5.1 Procedures for NFT Minting and Trading in the Marketplace
  • Mint. User can mint their NFT from the project website or NFT marketplace. In gernal, NFT marketplace have friendly tools to help creators deploy and mint a new NFT.
@@
  • Ownership Transfer. When someone purchases an NFT, the ownership of that digital asset is transferred to their digital wallet. The transaction is recorded on the blockchain, providing an immutable record of ownership.
@@ 5.2.1 Marketplaces
Blur [24] is an innovative NFT marketplace that caters to professional traders. Launched in 2022, Blur quickly surpassed OpenSea in terms of volume due to its low fees, high speed, and unique features such as NFT floor depth charts. Blur is more than just a marketplace; it’s an ecosystem consisting of an NFT marketplace, aggregator [39], and lending platform [47]. The Blur marketplace provides advanced analytics, portfolio management features, and the ability to compare NFTs on different marketplaces. The marketplace’s main feature is that it allows for gasless transactions, which significantly reduces transaction costs for users [161]. To get started, users first connect their wallet and sign a gasless transaction. Then, they start listing and bidding on NFTs. The platform encourages users to upload or list NFTs to qualify for airdrops. Blur has its native token, which has been distributed to the community over several quarters.
@@
SuperRare [179] is a leading platform for buying, selling, and collecting digital art as NFTs. Launched in 2018, it focuses on high-quality, limited-edition digital artwork created by both emerging and established artists. One of SuperRare’s key features is its curation process, which ensures that top-notch artwork is featured on the platform. Artists must apply or be invited to join SuperRare, and each piece of art goes through a thorough review by the platform’s curation team. This strict curation helps maintain the integrity of the art on SuperRare, making it a trusted marketplace for collectors. The platform offers a vibrant marketplace where users can discover and acquire digital art in various styles, media, and themes. SuperRare supports different formats, including static images, animated GIFs, and even interactive 3D experiences, allowing artists to explore new ways of expressing themselves digitally. A unique aspect of SuperRare is its focus on provenance and royalties. Every time an artwork is resold, a portion of the sale proceeds automatically goes to the original creator, ensuring that artists continue to benefit from the increasing value of their work over time.
@@
Magic Eden [124] is a decentralized marketplace based on the Solana blockchain, focusing on buying, selling, and trading NFTs. Launched to meet the growing demand for NFTs and digital collectibles, Magic Eden offers a range of features designed to provide a smooth experience for users within the Solana ecosystem. At its core, Magic Eden is a marketplace where users can discover, buy, and sell various NFTs, including digital art, collectible cards, virtual items, and more. The platform has an intuitive interface that makes it easy to browse and explore different collections, helping users find unique and valuable NFTs that match their interests. One of Magic Eden’s main advantages is its integration with the Solana, known for its high performance, scalability, and low transaction fees. By leveraging Solana’s capabilities, Magic Eden provides fast and cost-effective transactions, allowing users to buy, sell, and trade NFTs with minimal friction. Additionally, Magic Eden offers features like staking, governance, launchpads, and community engagement. Users can stake SOL tokens to earn rewards, participate in governance decisions, and contribute to the platform’s growth.
Table 6 shows the main features for the NFT marketplace. The Figure 7 provides the NFT marketplace share from Dec 2021 to Sep 2024.
NFT Pricing. NFTs represent a unique class of digital assets, often associated with digital art, collectibles, and virtual real estate. Unlike fungible cryptocurrencies, each NFT is distinct, often carrying unique metadata that determines its individual value. Moreover, there is much research on NFT prices, as shown in Table 7. Understanding the price determinants of NFTs involves different factors that includes textual and visual data, market dynamics, social media influence, and broader financial trends. Textual descriptions and image data within NFT collections can help explain price variations among individual NFTs; however, features extracted from these data do not generalize well to new, unseen collections [237]. This indicates that, while certain stylistic or descriptive elements might boost an NFT’s value within a specific collection, these factors do not necessarily apply across different collections. Interestingly, NFTs described with negative sentiment tend to fetch higher prices than those with positive sentiment [146]. This could be attributed to the perceived uniqueness or controversial nature of such descriptions, which might attract more attention and demand.
@@
Title Data Source Time Period Topic AI Used AI Model Analysis Data Analysis
@@
NFT markets are relatively immature and often inefficient [52]. For instance, the pricing of virtual real estate in Decentraland does not always reflect all available information, though a general rise in value has been observed over time. This inefficiency suggests opportunities for speculative gains but also underscores the risks involved. NFTs have been found to offer significant diversification benefits within traditional asset-based portfolios [107]. During stable market times, NFTs can act as transmitters of systemic risk, but in the stressful periods, they tend to absorb risk spillovers [9]. This dual role highlights their potential as both a hedge and a diversification tool in investment portfolios. The popularity of NFTs are heavily influenced by the performance of major cryptocurrencies like BTC and Ether [158]. Increased prices could drive hype and attention towards NFTs, as evidenced by co-movement patterns observed through wavelet coherence analysis [53]. Furthermore, Bitcoin returns have been shown to predict NFT growth in popularity, as measured by Google search queries.
The key factors contributing to NFT value have been identified: rarity, reputation, authenticity, significance, utility, interoperability, demand, provenance, royalties, and hype [115]. These aspects encapsulate both intrinsic qualities of NFTs and external market perceptions that drive their valuation. During extreme market conditions, various economic indicators such as equity and gold market uncertainties, business conditions, and geopolitical risks significantly predict returns in specific NFT markets [193]. This indicates that broader economic factors can play a crucial role in determining NFT prices. The valuation of NFTs is influenced by a complex interplay of textual and visual data, social media presence, market dynamics, financial contagion, and broader economic conditions. These factors collectively shape the perceived value of NFTs.
@@ 5.3 NFT Security
The rapid proliferation of NFTs market is facing with challenges in cybersecurity that pose risks to artists and collectors alike, including both human and technological issues. Figure 8 shows most of the security issues for NFT trading.
@@
Technological vulnerabilities are common because of the complex nature of smart contracts and the intricate design of NFTs. Smart contracts, which are self-executing contracts with the terms of the agreement directly written into code [101], are foundation to the operation of NFTs. However, they are not immune to security flaws. Identity verification issues might allow malicious attackers to impersonate legitimate users and lead to unauthorized transactions. The verifiability of NFT contracts is another critical issue. NFT contract is authentic and has not been tampered with is essential for maintaining trust in the system. Smart contract vulnerabilities, such as reentrancy, transaction order dependence, integer overflows, and unhandled exceptions, all highlight the security challenges between NFTs and broader DeFi ecosystems [26, 68, 76, 92, 97, 122].
@@
In addition, NFT market suffers from a lack of transparency [38], particularly in off-chain trades where transactions occur outside the blockchain. This opacity can facilitate fraudulent activities and make it difficult to verify the authenticity and ownership history of an NFT. Fairness in bidding processes is another concern, with practices such as shill bidding and bid shielding undermining the integrity of auctions. Royalty and marketplace fee evasion pose financial risks to artists and platforms, while cross-platform issues and non-enforcement of contractual terms can further exacerbate financial risks.
Security issues in Web2, such as Denial of Service (DoS) attacks [139], also persist in the NFT ecosystem. With substantial funds flowing into DeFi applications, the allure of scams and cyber attacks has grown. Cryptoeconomic strategies, such as transaction reordering [100, 118], flash loan abuse [164], arbitrage opportunities [44], and pump-and-dump schemes [221, 98], have been adapted to exploit vulnerabilities in the NFT market.
@@
Wash Trading. The emergence of NFTs has ushered in a new era of digital asset markets, characterized by unprecedented innovation and rapid growth. However, amidst the excitement surrounding NFTs lies a darker underbelly: the prevalence of wash trading (cf. Figure 9). Wash trading, a deceptive practice involving the simultaneous buying and selling of assets to create artificial trading volume, has raised significant concerns within the NFT ecosystem. Recent studies challenged previous estimates, suggesting that wash trading may be less common than initially thought [194]. Nonetheless, findings still indicate alarming levels of wash trading activity, with some estimates it attributing up to 25% of total trading volume to this illicit practice [189].
@@
Identifying and combating wash trading in NFT markets presents unique challenges. Innovative solutions are being proposed, such as algorithms to detect wash trading at different levels, from transaction to collector [182]. These efforts aim to provide regulators and investors with tools to navigate the complexities of the NFT landscape and mitigate the risks associated with market manipulation. Moreover, the motivations behind wash trading are multifaceted. While some engage in wash trading to artificially inflate NFT prices, others exploit token reward systems offered by marketplaces  [113]. The profitability of wash trading is evident, with some perpetrators avoiding significant losses and even profiting from reselling NFTs to unsuspecting buyers.As the NFT market continues to evolve, the need for robust regulation becomes increasingly apparent. Measures must be implemented to prevent market abuse, instill investor confidence, and safeguard the integrity of digital asset markets. Novel visualization tools, like NFTDisk [210], offer investors a means to identify wash trading activities and make informed decisions in an increasingly complex market landscape. Table 8 provides the summary of the identified literature for NFT Wash Trading.
@@
Unveiling the Paradox of NFT Prosperity [83] OpenSea, CryptoPunks, LooksRare, X2Y2, Blur ERC-721 ERC-1155 Wash trading analysis ACM
@@
Identifying Washtrading Cases in NFT Sales Networks [182] Nifty Gateway Wash trading detection IEEE Transactions
@@
Beyond the Surface: Advanced Wash Trading Detection in NFT Markets [189] Blur, OpenSea, X2Y2 ERC-721 LooksRare Wash trading analysis arVix
@@
NFTDisk: Visual Detection of Wash Trading in NFT Markets [210] OpenSea, LooksRare ERC-721 Wash trading detection ACM